THE A-X AND C - X TRANSITIONS IN Kr.NO

\underline{TIMOTHY G. WRIGHT}, J. M. DYKE, P. MACK and D. M. SMITH, CHEMISTRY DEPART-MENT, UNIVERSITY OF SOUTHAMPTON, HIGHFIELD, SOUTHAMPTON SO17 1BJ, U. K..

The Kr.NO complex has been studied using resonance-enhanced multiphoton ionization spectroscopy. More-detailed spectra from those reported previously have been obtained. The A-X transition was reported last year (Abstract RE05). A brief summary of the information obtained, and the non-Rydberg behaviour of these molecules will be presented.